OpenStudy (anonymous):

10 practice questions!! need help! what is the value of the expression? 12+6*2/18-3*5

OpenStudy (skullpatrol):

Order of operations: 1. Do all multiplications and divisions in order from left to right. 2. Do all additions and subtractions in order from left to right.

sammixboo (sammixboo):

Does your problem look like this? \(\rm \Large \frac{12~+~6\times 2}{18~-~3\times 5}\)

OpenStudy (anonymous):

yes it does. i didnt know how to make it look like that

sammixboo (sammixboo):

Ah okay So make sure you follow the Order of Operations P - Parenthesis E - Exponents M & D - Multiplication & Division A & S - Adding and Subtracting As we see there is no parenthesis nor exponents, so we move on the Multiplying and Dividing \(\rm \Large \frac{12~+~\color{red}{6\times 2}}{18~-~\color{red}{3\times 5}}\) So what is \(\rm \color{red}{6\times 2}\) and \(\rm \color{red}{3\times 5}\)

OpenStudy (anonymous):

6*2= 12 and 3*5=15

sammixboo (sammixboo):

And like @skullpatrol said, make sure you do the multiplication and subtraction left from right, and same with addition and subtraction.

OpenStudy (anonymous):

ok so it would be 12+12 divided by 18-15

sammixboo (sammixboo):

Good! Now we plug in \(12\) where \(6\times 2\) was, and \(15\) where \(3 \times 5\) was in the equation \(\rm \Large \frac{12~+~\color{red}{6\times 2}}{18~-~\color{red}{3\times 5}}\) \(\rm \Large \frac{12~+~\color{red}{12}}{18~-~\color{red}{15}}\) Now we must do the addition and subtraction. What is \(12~+~12\) and \(18~-~15\)?

OpenStudy (anonymous):

24 divided by 3 which equals 7

OpenStudy (anonymous):

no i mean 8

sammixboo (sammixboo):

Good job! That is your answer!

sammixboo (sammixboo):

Yeah * lol xD

sammixboo (sammixboo):

If you need any more help tag me \(\rm @sammixboo\)

sammixboo (sammixboo):

\(\LaTeX~Check\)

OpenStudy (anonymous):

ok thank you